### v1

# DSAI: Unbiased and Interpretable Latent Feature Extraction for Data-Centric AI

Abstract
Large language models (LLMs) often struggle to objectively identify latent characteristics in large datasets due to their reliance on pre-trained knowledge rather than actual data patterns. To address this data grounding issue, we propose Data Scientist AI (DSAI), a framework that enables unbiased and interpretable feature extraction through a multi-stage pipeline with quantifiable prominence metrics for evaluating extracted features. On synthetic datasets with known ground-truth features, DSAI demonstrates high recall in identifying expert-defined features while faithfully reflecting the underlying data. Applications on real-world datasets illustrate the framework’s practical utility in uncovering meaningful patterns with minimal expert oversight, supporting use cases such as interpretable classification  [^1] The title of our paper is chosen from multiple candidates based on DSAI-generated criteria..
DSAI: Unbiased and Interpretable Latent Feature Extraction for Data-Centric AI
Hyowon Cho1, thanks: Work done while interning at NAVER AI Platform, Soonwon Ka 2, Daechul Park2, Jaewook Kang2, Minjoon Seo1,thanks: Corresponding authors., Bokyung Son2,00footnotemark: 0 1KAIST AI, 2 NAVER AI Platform, 1{hyyoka, minjoon} @kaist.ac.kr 2{soonwon.ka, daechul.park, jaewook.kang, bo.son} @navercorp.com
1 Introduction
The ability to analyze large-scale datasets is a cornerstone of deriving actionable business insights. Traditionally, this task has been managed by human data scientists, but it faces several key challenges: (1) the large volume of data makes it difficult to review all information comprehensively, (2) human analysis can often include subjective bias, and (3) collaboration with domain experts is often required, leading to high operational costs.
Large language models (LLMs) have emerged as powerful tools for identifying patterns within massive datasets, leveraging their ability to process and generate language in context Touvron et al. (2023); Dubey et al. (2024); Achiam et al. (2023); OpenAI (2024); Lam et al. (2024); Wan et al. (2024). However, their application to data analysis is limited by critical shortcomings. First, LLMs often struggle to identify latent characteristic patterns in large datasets due to inherent data grounding issues, where outputs rely on pre-trained knowledge rather than the specific nuances of the input data Kossen et al. (2024); Kenthapadi et al. (2024); Wu et al. (2024). Second, the difficulty in verifying LLM-generated responses and the lack of quantitative evaluation methods require expert oversight, which can be prohibitively expensive at scale.
Figure 1: DSAI automates the extraction of latent features like syntax, domain, and clarity
To address these limitations, we propose Data Scientist AI (DSAI), a framework that systematically applies LLMs to extract and refine latent features from data. Unlike direct feature extraction approaches, DSAI adopts a bottom-up approach, starting with detailed analysis of individual data points, aggregating their characteristics, and deriving actionable features. The process is guided by defined perspectives which provide LLMs with a consistent framework for interpreting data points while minimizing subjective bias.
The DSAI pipeline operates in five stages: #1 Perspective Generation identifies data-driven perspectives from a small subset of data. #2 Perspective-Value Matching assigns values to individual data points by evaluating them against these perspectives. #3 Clustering groups values with shared characteristics to reduce redundancy. #4 Verbalization converts extracted features into a compact criterion form. #5 Prominence-based Selection determines which features to use based on a prominence intensity metric that quantifies the discriminative power of each extracted feature. Throughout the process, LLMs remain task-agnostic to minimize bias, ensuring that their outputs are grounded in data rather than domain-specific assumptions.
To validate our framework, we demonstrate its effectiveness on datasets specifically curated for our experiments, which include research titles Wang et al. (2018) and slogans Jin et al. (2023). We then showcase DSAI’s practical value using real-world datasets: news headlines with CTR Wu et al. (2020), spam detection Kim (2016), and Reddit comments with community engagement metrics Magnan (2019). These experiments highlight DSAI’s ability to provide actionable insights for industries focused on content optimization, user engagement, and moderation.
Our main contributions are as follows: (1) Minimizing bias by ensuring LLMs identify latent characteristics within large datasets; (2) Introducing a quantitative metric for feature prominence scoring, which serves as a proxy for measuring discriminative power of each feature; (3) Improving interpretability through feature-to-source traceability. (4) Enabling thorough examination of large datasets while minimizing human labor through systematic guidance of LLMs;
2 Related Works
2.1 Latent Feature Extraction with LLM
Recent advancements in LLMs have demonstrated their effectiveness in extracting latent features, particularly in identifying perspectives and matching values in data Peng et al. (2023). Studies using LLM-based clustering techniques have shown promising results in extracting high-level conceptsLam et al. (2024); Wan et al. (2024); Viswanathan et al. (2024); Pham et al. (2024), demonstrating their utility for analyzing large datasets Wang et al. (2023); Kwon et al. (2023). Research has also shown that decomposing complex tasks into multiple stages or aspects enhances performance Saha et al. (2023); Liu et al. (2024). Our framework combines perspective generation, multi-stage feature construction, and clustering using LLMs to conduct comprehensive latent feature extraction.
2.2 Bias of LLMs
LLMs face challenges in adapting to new patterns due to pre-existing knowledge biases. Kossen et al. (2024) show that in-context learning (ICL) struggles to overcome these biases even with explicit prompts or many-shot examples. Using external knowledge bases also has limited effectiveness in reducing hallucinations and reliance on internal biases Kenthapadi et al. (2024); Lee et al. (2023). Advanced LLMs fail to align with provided context in about 40% of predictions when the given context conflicts with their prior knowledge Wu et al. (2024). These findings emphasize the need for robust techniques to mitigate bias and enhance grounding, especially for applications where data-driven conclusions are crucial.
3 Challenges in LLM-Driven Data Analysis
  Reference Criteria Comparison Criteria Recall (%)
Slogan Title
  FlippedPosData PosData 89.5% 94.1%
FlippedMixedData MixedData 81.8% 89.5%
NoData PosData 100.0% 95.6%
MixedData 100.0% 92.3%
FlippedPosData 90.0% 94.1%
FlippedMixedData 96.4% 85.2%
Expert NoData 88.9% 83.3%
PosData 66.7% 54.2%
MixedData 50.0% 75.0%
FlippedPosData 94.4% 45.8%
FlippedMixedData 77.8% 83.3%
DSAI (Thres: [0]) 100.0% 83.3%
DSAI (Thres: [0.348]) 88.9% 83.3%
DSAI (Thres: [0.692]) 77.8% 75.0%
Table 1: Recall of features derived from different approaches across slogans and titles.
This section demonstrates that LLMs rely more on pre-trained knowledge than data-specific features, showing a tendency to depend on pre-trained knowledge when input data is absent or labels are flipped, and generating vague features when task-specific context is removed.
3.1 Setting
Dataset Annotation and Sampling
Our experimental task aims to identify the characteristics of high-quality text as defined by expert criteria Nair and Gibbert (2016); Tullu (2019); Kohli et al. (2007); Padrakali and Chitra Chellam (2017). We use two distinct datasets: (1) commercial slogans and (2) academic research titles, representing different text genres (Appendix A).
To construct the high-quality sample set, we used LLMs to evaluate each entry in the 3,000-entry dataset according to expert criteria. Binary scores were assigned for each criterion, and final ratings were aggregated. The top 600 samples formed the analysis group ("positive"), and the bottom 600 formed the control group ("negative").222”Positive”/”negative” indicates the presence/absence of specific features rather than overall quality. See Appendix C for detailed annotation procedures.
Model
We used GPT-4o OpenAI (2024) for its advanced capabilities in nuanced text understanding and annotation tasks Tan et al. (2024).
3.2 Data Grounding Issues
We explore directly extracting latent features using LLMs with two different types of input data:
  • PosData: Uses only positive (top-scored) samples to identify characteristics of high-quality samples.
  • MixedData: Uses both positive (top-scored) and negative (bottom-scored) samples to capture a broader range of features.
Both approaches have generated features that substantially overlap with expert-defined criteria (Table 1). However, since our datasets come from popular domains, it is unclear whether the LLM-generated features are truly derived from the provided data or from the model’s pre-existing domain knowledge. To test this, we pose several key questions and conduct experiments to evaluate them.
(a) Does the LLM Adapt to Input Data?
We tested whether the LLM uses input data to adapt its feature generation. To evaluate this, we flip the labels and observe if the generated features change accordingly:
  • FlippedPosData: Negative examples are labeled as positive.
  • FlippedMixedData: Positive and negative labels are reversed.
As shown in Table 1, features generated from PosData and FlippedPosData exhibited minimal changes, with significant overlap also observed between MixedData and FlippedMixedData. This indicates that LLMs rely more on pre-existing knowledge than on input data for feature generation.
(b) Is Pre-existing Knowledge the Primary Source of Generated Features?
To determine if features are primarily sourced from internal knowledge, we generated features without any input data:
  • NoData: The LLM generates features solely based on its pre-trained knowledge.
Table 1 shows that NoData achieved high recall on expert-defined criteria and significant overlap with PosData, MixedData, and flipped datasets. These findings confirm that pre-existing knowledge, rather than data-specific information, drives LLM-generated features.
(a) DSAI shows superior grounding over direct feature generation methods, with no DP score below 0.5. When prominence exceeds 0.2, all scores remain above 0.6.
(b) The graph shows that higher prominence leads to higher DP scores for DSAI features.
Figure 2: DP scores for direct feature generation and DSAI methods.
(c) Are the Generated Features Truly Reflective of Positive Data’s Latent Characteristics?
We computed the discriminative power (DP) of features to assess whether they reflect the latent characteristics of positive data (details in Section 5.2). Features with DP values $\leq 0.5$ fail to differentiate between positive and negative traits, while values $>0.6$ indicate strong discrimination.
As shown in Figure 2(a), several features from PosData and MixedData exhibited low DP scores, some below 0.5. These results highlight the limitations of LLMs in generating features that truly represent positive data’s latent traits.
(d) Does Removing Task Context Improve Data Grounding?
To test if removing task-specific context reduces bias and improves data grounding, we used:
  • NoContext: The LLM receives only unlabeled data groups and is unaware of the goal of identifying high-quality text.
In this setting, the LLM defaulted to generating vague and generic descriptors (Appendix D). This suggests that task context removal alone does not improve data grounding and highlights the need for structured guidance during feature extraction.
Figure 3: Overview of the DSAI pipeline: Perspectives are first generated to guide analysis (#1), then used to match values to data points (#2). These values are clustered to reduce redundancy (#3), verbalized into concise criteria (#4), and prioritized based on their prominence (#5).
4 Data Scientist AI
DSAI is a 5-stage framework for automated latent feature generation, leveraging LLM capabilities with structured processes (Figure 3). Below, we outline the DSAI pipeline and its five stages.
#1 Perspective Generation
LLMs can extract diverse attributes from text Peng et al. (2023). Using a subset of positive and negative samples, this stage generates distinct attributes while minimizing reliance on domain bias by excluding task-specific context. Each perspective includes a name, evaluation criterion, analysis process, and example (Figure 3 (1)). This structure provides a consistent framework for dataset interpretation. After generating perspective candidates, duplicates are removed.
#2 Perspective-Value Matching
This stage assigns values to each data point guided by the generated perspectives on the previous stage.
#3 Value Clustering
This stage groups similar values within each perspective to create interpretable feature sets with reduced redundancy. Based on LLM clustering literature Wang et al. (2023); Wan et al. (2024); Lam et al. (2024), the process consists of two steps: generating representative cluster labels, and assigning values to appropriate clusters (Figure 3 (3)).
#4 Verbalization
This stage transforms perspective-label pairs into verbal criteria. For each pair, we compute $P(\textrm{positive}|D_{p,l})$, the proportion of positive examples in dataset $D_{p,l}$ corresponding to the perspective-label pair $(p,l)$. Using this, we calculate a directional score $12\times P(\textrm{positive}|D_{p,l})-$ 1, which determines how the pair should be verbalized. Pairs with positive directional scores (>0) are directly verbalized as features describing positive data. Pairs with negative directional scores (<0) are transformed into "avoid" statements, which indirectly characterize positive data by specifying features to avoid. [^3] For instance, if (clarity, low) receives a negative directional score, it is verbalized as ”Avoid sentences with low clarity.” This dual transformation approach ensures coverage of both desired and undesired traits.
#5 Prominence-based Selection
Finally, DSAI employs prominence intensity as the feature selection metric, defined as the absolute value of directional score $\|2*P(\textrm{positive}|D_{p,l})-1\|$. This metric indicates how decisively a perspective-label pair characterizes either positive or negative samples–the closer to 1, the stronger the association. Unlike our baseline approaches (Section 3.2) that merely output a list of features without any relative importance measures, our prominence intensity metric provides a quantitative way to assess and compare the discriminative power of each feature. Users can set a prominence intensity threshold to match their specific needs, balancing between stronger predictive patterns (higher threshold) and broader coverage of the dataset’s latent characteristics (lower threshold).
Traceability and Transparency
The feature generation process maintains the link of each feature to its perspective-label combination, the original data samples that contributed to its creation, and the quantitative prominence score that justified its selection. Such traceability enables users to verify the legitimacy of extracted features by examining their source data and the complete derivation process.
5 Validation of Methodology Using Expert-Driven Annotation Dataset
This section validates our methodology through experiments on various datasets, focusing on three key aspects: recall of expert-defined criteria (§5.1), discriminative power of generated criteria (§5.2), and reliability of pipeline modules (§5.3).
5.1 Recall of Expert-Defined Criteria
We evaluate DSAI’s ability to reproduce expert-defined criteria by applying it to slogan and research title datasets annotated as described in Section 3.1.
Criteria Generation
We generated criteria through our pipeline and retained those with $D_{p,l}>6$ and positive prominence intensity scores. This yielded 235 criteria for slogans and 198 criteria for research titles.
Human Feature Matching
For recall evaluation, one annotator initially performed loose matching of generated criteria, which was then validated through majority voting among three annotators.
Recall
Our methodology showed strong performance in reproducing expert-defined criteria, even at high prominence intensity thresholds. All 9 expert criteria for slogans were captured at a threshold of 0.348, and 83% recall (10/12) was achieved for research titles at a threshold as high as 0.692 (Table 1). While PosData and MixedData also showed decent coverage, their results relied on LLM’s pre-existing knowledge as discussed in Section 3.2. In contrast, our approach, by design, minimizes such potential bias by withholding task-specific context, while still achieving comparable or better recall rates.
Recall Dynamics across Various Thresholds
The adjustable prominence intensity thresholds allow users to tailor their analyses by balancing discriminative power and coverage. Additional analyses on recall dynamics across various thresholds are provided in Appendix H.
5.2 Discriminative Power (DP)
Figure 4: Example of interpretable spam classification: The figure shows how feature prominence guides criteria selection, with high-prominence criteria improving spam classification performance.
To validate our pipeline’s effectiveness, we examined whether criteria with higher prominence intensity scores indeed showed stronger discriminative power in practice. We divided DSAI-generated criteria into five buckets based on prominence intensity scores to cover the full range. From each bucket, 10 features were sampled and used to annotate datasets for feature conformity ("conform"/"not-conform") where applicable. Non-applicable cases444e.g., features referencing business models in unrelated data. were excluded from calculations.
We define DP score as follows:
$DP\text{ }Score=\left\{\begin{aligned} &P(\textrm{positive}|\textrm{conform})\quad\text{if }\\ &\quad\quad\quad\quad\quad\textrm{directional\_score}>0,\\ &P(\textrm{negative}|\textrm{notConform})\quad\text{elif }\\ &\quad\quad\quad\quad\quad\textrm{directional\_score}<0.\end{aligned}\right.$
DP scores above 0.5 indicate effective discrimination of positive examples, while scores below 0.5 suggest potential errors in the LLM-based pipeline (either in direct or "avoid" form). Criteria scoring between 0.5 and 0.6 were deemed too generic to provide meaningful discrimination.
As shown in Figure 2(b), criteria with higher prominence scores demonstrated stronger discriminative power, with all DSAI-generated criteria scoring above 0.5. This validates not only our pipeline’s reliability but also the effectiveness of the prominence intensity metric in identifying meaningful discriminative criteria. Through this metric, users can quantitatively assess which features are more reliable indicators of positive samples, rather than treating all extracted features as equally important.
5.3 Reliability of Pipeline Operations
Due to the inherent uncertainty in LLM outputs, we evaluated the reliability of pipeline stages that rely on LLM operations by having the LLM verify its own mappings between datapoints and features. We assessed three LLM-dependent stages by asking the LLM to confirm the reliability of its outputs. [^5] Stage #1 (Perspective Generation) was not evaluated as perspective generation allows for more diverse outputs. Stage #5 (Prominence-based Selection) was excluded as the results are deterministic. Our verification process showed high consistency rates of >98%,  94%, and 98% for stages #2, #3, and #4 respectively (details in Appendix J).
6 Real-World Application with Quantitative Datasets
We applied DSAI to three real-world user feedback datasets critical for business insights Luo et al. (2022): (1) MIND Wu et al. (2020), analyzing engagement features in news headlines with high CTR as the positive group; (2) spam detection Kim (2016), identifying patterns in spam messages as the positive group; and (3) Reddit Magnan (2019), exploring interaction-promoting linguistic features in highly upvoted comments as the positive group.
Prominence Distribution and Samples
The datasets show distinct patterns, highlighting the need for dataset-specific thresholds to balance discriminative power, recall, and domain priorities (e.g., minimizing spam errors or maximizing engagement). In addition to general characteristics, DSAI-generated features notably capture more specific and fine-grained nuances across various subsets of a given dataset, often overlooked in LLM-based analyses. Detailed analysis is provided in Appendix L.
Potential for Downstream Tasks
Our methodology provides latent characteristics of data in linguistic form, enabling diverse applications such as style transfer, annotation guideline generation, and classification tasks. To demonstrate the practical utility of DSAI-generated criteria, we conducted a toy spam classification experiment using 20 spam and 20 ham samples. Using the five criteria with the highest prominence intensity scores led to effective classification performance, while using the five criteria with the lowest prominence resulted in poor performance (See Figure 4). This illustrates DSAI’s potential to establish interpretability along with feature credibility through prominence scores, making it suitable for automatic application in various downstream tasks including classification.
7 Conclusion
In this paper, we proposed DSAI, a faithful data-driven feature extraction framework that ensures LLMs identify latent characteristics from data without relying on their domain-related biases. DSAI automates thorough examination of large datasets while minimizing human labor and enhancing interpretability through source-to-feature traceability. Through empirical validation, we confirmed its capability to extract meaningful features, suggesting its potential for applications requiring interpretable and efficient feature extraction.
Appendix A Details of Datasets
To verify the applicability and effectiveness of our methodology, we utilized datasets from diverse industries, featuring a wide range of attributes and characteristics. Two datasets (slogan and research title) were used to validate our approach, while the other datasets were employed to demonstrate DSAI’s practical applications across various domains.
Slogan Dataset
We utilized the Slogan Generation Dataset Jin et al. (2023), a publicly available resource containing 5,014 validation examples, 1,322 test examples, and 251 human-annotated examples rated for "catchiness." Each entry includes a company name, a description of the company or product, the slogan, and its industry category. For example, the dataset features entries like Bikini Village Inc., categorized under "Apparel & Fashion," with the slogan "Your Brand Destination For Swimwear" and a description emphasizing its focus on trendy swimwear and beachwear. The dataset primarily features lesser-known companies, reducing bias from brand recognition and enabling analysis to focus on the intrinsic qualities of slogans across various industries.
Research Title Dataset
For analyzing research titles, we utilized a dataset of 10,874 research paper title and abstract pairs from previous work Wang et al. (2018). This dataset was preprocessed and divided into training (80%), validation (10%), and testing (10%) sets. On average, each title contains 9 words, and abstracts have 116 words.
MIND Dataset
The MIND dataset, created for news recommendation tasks, is a large-scale benchmark dataset sourced from Microsoft News Wu et al. (2020). It includes data from 1 million users and over 160,000 English news articles, each containing a title, abstract, and body text. We calculated click-through rates (CTR) based on the provided click and rejection logs for each article.
Spam Dataset
The SMS Spam Collection is a widely used dataset for spam detection research, containing 5,574 SMS messages in English, labeled as either "ham" (legitimate) or "spam" Kim (2016). This dataset was compiled from several public sources, including Grumbletext (425 manually extracted spam messages), the NUS SMS Corpus (3,375 legitimate messages from Singaporean students), Caroline Tag’s Thesis (450 ham messages), and the SMS Spam Corpus v.0.1 Big (1,002 ham messages and 322 spam messages).
Reddit Dataset
The Reddit dataset is a subset of a larger dataset from May 2019, containing 25,000 comments from each of the 40 most active subreddits Magnan (2019). Comments that were removed, deleted, or shorter than 4 tokens were excluded. The dataset includes features such as the subreddit where the comment was posted (categorical), the comment content (string), a binary controversiality metric, and a score indicating net upvotes minus downvotes.
Appendix B Expert-Defined Criteria
Previous research has devoted considerable effort into identifying the characteristics of "good content" in specific domains. Building upon this foundation, we conducted a preliminary investigation to establish expert-defined criteria for both slogans and research titles.
Slogan
To establish criteria for effective slogans, we synthesized insights from two foundational studies. The first study emphasized how slogans enhance brand identity, equity, and differentiation, providing guidelines such as ensuring long-term relevance, clear brand positioning, consistent brand association, creative depth for better recall, and minimal false recall Kohli et al. (2007). The second study outlined core characteristics of effective slogans, highlighting simplicity, pleasantness, compactness, persuasiveness, memorability, uniqueness, and honesty Padrakali and Chitra Chellam (2017). After thoroughly reviewing the two studies, we synthesized their insights into a unified set of expert-defined criteria as in Figure 11.
Title
For research titles, we similarly synthesized insights from two key studies Nair and Gibbert (2016); Tullu (2019). These studies offer comprehensive guidance on crafting impactful titles that align with academic expectations. They both emphasize that effective research titles should be clear, concise, and informative, accurately representing the study while capturing readers’ attention. Key recommendations include placing essential keywords early for indexing purposes, avoiding ambiguity and misleading content, and maintaining professional language. The studies stress the importance of balanced structure and linguistic clarity, with descriptive titles preferred over declarative or interrogative formats. We integrated these recommendations into a comprehensive set of expert-defined criteria as in Figure 11.
Appendix C LLM Annotation with Expert-Defined Criteria
This section describes our methodology employed for data annotation using expert-defined criteria and evaluates the effectiveness of each criterion through distribution analysis.
C.1 Method
We established a gold standard by manually annotating 10–20 samples per dataset and using expert-defined criteria. After optimizing prompts to maximize alignment with this manual annotation, we used the most aligned prompt to annotate 3,000 data points with GPT-4o, classifying each as "Conform" or "Not conform". The conformity score for each data point was calculated as the total number of conforming criteria. Based on these scores, we selected the top and bottom 600 samples as positive (high-quality text) and negative (low-quality text) samples respectively.
This annotation process proved cost-effective compared to human annotation, with an average cost of \$3.5 USD per 3,000 annotations.
  Criterion Top Conform (%) Bottom Conform (%)
Direct/Straight-forward 99 61
Consise but not too simple 100 0
Pleasant to hear 84 1
Includes a sales idea 98 4
No exaggeration 99 91
Future-oriented 81 1
Clear positioning 93 2
Highlight the brand’s unique traits 99 0
Include the brand name 99 30
Table 2: Ratio of Conform in Top and Bottom for Slogan expert-defined criteria.
C.2 Ensuring Distinction Between Positive and Negative Groups
Given that some expert criteria may lack sufficient discriminative power despite their theoretical importance, we conducted an analysis comparing the ratios of "Conform" for each criterion between top- and bottom-ranked samples. Note that a truly discriminative criterion should show high conformity in the top group and low conformity in the bottom group.
  Checklist Item Top Conform (%) Bottom Conform (%)
Simple format 100 3
Direct 100 1
Informative and specific 100 0
Functional (with scientific keywords) 100 3
Concise and precise 100 0
Include the main theme 100 0
Not too long or too short 100 3
Avoid whimsical words 100 60
Avoid jargon 100 44
Mention place/sample size if valuable 100 72
Important terms at the beginning 100 11
Descriptive titles preferred 100 9
Table 3: Ratio of Conform in Top and Bottom for Title expert-defined criteria.
Slogan (Top vs. Bottom)
Table 2 demonstrate significant conformity ratio disparities between the top and bottom groups for most criteria. Several criteria demonstrate maximal distinction (e.g., "Concise but not too simple" (100% vs. 0%)), and some show moderate but still meaningful differentiation (e.g., "Include the brand name" (99% vs. 30%)). "No exaggeration" (99% vs. 91%) and "Direct/Straight-forward" (99% vs. 61%) have smaller gaps.
Title (Top vs. Bottom)
Table 3 shows even stronger distinction between the top and bottom groups. Multiple criteria (e.g., "Simple format" "Direct", "Concise and precise") achieve 100% conformity in the top group compared to as low as 0–3% conformity in the bottom group, indicating their high predictive value for title quality. The large differences across all criteria confirm their strong discriminative power.
Appendix D Direct Latent Feature Extraction using LLMs
To evaluate whether the features directly generated by large language models (LLMs) are grounded in the input data or influenced by pre-existing knowledge, we conducted a series of experiments. These experiments were designed to investigate the effects of input data, label variations, and task context on feature generation (Section 3.2).
In these experiments, we ensured that all available data were utilized to maximize the representativeness of the input. For instance, we used all 600 positive samples for PosData and the complete set of 1200 samples (both positive and negative) for MixedData. This comprehensive use of data aimed to minimize bias and ensure that the features generated by the LLM were robustly grounded in the provided input.
To further address potential data-grounding issues in the direct feature generation process, we explicitly included instructions in the prompts emphasizing that the LLM must consider the input data when generating criteria. By reinforcing the importance of grounding, we sought to mitigate the risk of features being overly influenced by the model’s pre-existing knowledge, while the results revealed that the generated features were still significantly influenced by the LLM’s pre-existing knowledge (See Results in Section 3.2).
The specific prompts employed in each experimental setup are detailed in Figure 13, which provides a comprehensive outline of the instructions given to the LLM. The results of the experiments are presented in Figures 14 to 22, offering insights into the effectiveness of the approach and highlighting the model’s performance across various configurations.
Appendix E Implementation Details and Cost Analysis
In this section, we provide the prompts used in each stage of the DSAI pipeline along with corresponding outputs.
Details of Each Inference
In Stage #1 Perspective Generation, we generated 50 perspectives at per forwarding step across three steps, iteratively concatenating each step’s output with the few-shot example in the prompt for the next step. For Stage #2 Perspective-Value Matching, each forwarding step processed one sentence and three perspectives as input. Stage #3 Perspective-Oriented Value Clustering used all generated values per perspective in order to effectively cluster shared characteristics. In Stage #4 Verbalization, we transformed each perspective-label pair into a compact sentence (criterion). The prompts for all stages are in Figures 23 to 27.
Cost Analysis
The total cost of processing 10 perspectives and 100 sentences is \$2.43746, broken down as follows:
In Step 1: Perspective Generation, generating 10 perspectives costs \$0.0304. Each perspective costs \$0.00304. With a total of 2,100 input tokens and an average of 82.90 output tokens per perspective, the cost is calculated as:
$10\times 0.00304=0.0304$
In Step 2: Perspective-Value Matching, each sentence requires $\lceil 10/3\rceil=4$ inferences, as each inference processes 3 perspectives. For 100 sentences, this results in:
[…]
With a cost per inference of \$0.00496, the total cost for this step is:
$400\times 0.00496=1.984$
On average, each inference involves 2,206 input tokens and 440 output tokens.
In Step 3: Perspective Value Clustering, each perspective incurs a clustering cost of \$0.0087125. This includes two components:
  • Label Generation, costing \$0.00579625 for 597 input tokens and approximately 1,010 output tokens.
  • Value Matching, costing \$0.00291625 for 269 input tokens and 516 output tokens.
The total clustering cost per perspective is:
$0.00579625+0.00291625=0.0087125$
For 10 perspectives, the total clustering cost is:
$10\times 0.0087125=0.087125$
In Step 4: Verbalization, it is assumed that the 10 perspectives cluster into 5 groups each, resulting in 50 perspective-label pairs. Verbalizing each pair costs \$0.0018, so the total cost for this step is:
$50\times 0.0018=0.09$
Summing all the costs, the total processing cost is:
$0.0304+1.984+0.087125+0.09=2.43746$
Appendix F Discriminative Power of Direct LLM- and DSAI-Generated Features
The accuracy with which generated features reflect the latent characteristics of the positive dataset depends on their alignment with the input data. To evaluate this alignment, we annotated the dataset using all generated features and assessed their discriminative power (DP) using two key metrics: $P(\textrm{positive}|\textrm{Conform})$ and $P(\textrm{negative}|\textrm{NotConform})$.
The metric $P(\textrm{positive}|\textrm{Conform})$ measures the proportion of positive data among instances labeled as "Conform" to a feature. This value indicates how effectively a feature captures the latent characteristics of the positive dataset. Scores below 0.5 suggest that the feature aligns more closely with negative data.
$P(\textrm{negative}|\textrm{NotConform})$ evaluates how successfully a feature identifies traits in the negative dataset. Scores above 0.5 indicate that the feature effectively captures the negative traits, while scores below 0.5 suggest misalignment, potentially indicating positive traits incorrectly marked as "Not Conform."
We used LLM labeling to filter out the "Not Applicable" cases (e.g., "technical terms in medical domains" for non-medical texts) before calculating DP scores.
This two-metric approach enables evaluation of features’ effectiveness in both capturing positive characteristics and identifying traits to avoid in negative data.
To simplify interpretation, in the following sections, we subtracted 0.5 from all DP scores. This adjustment highlights values below 0.5, making it easier to identify features that fail to align effectively with the intended dataset. Positive values after subtraction indicate alignment, while negative values signal misalignment.
F.1 Discriminative Power of LLM-Generated Features
Feature Slogan Title
PosData MixedData PosData MixedData
Feature 1 0.298535 0.061810 0.293475 0.214968
Feature 2 0.212195 0.193333 0.159765 0.075342
Feature 3 0.270186 0.295139 0.152727 0.049150
Feature 4 0.258410 0.031357 0.177354 0.169845
Feature 5 0.086310 0.069536 0.102517 0.160841
Feature 6 0.127063 0.120278 0.065525 0.170258
Feature 7 0.106327 0.227273 0.221347 0.191358
Feature 8 0.004625 0.128173 -0.116618 0.039169
Feature 9 0.333333 0.189394 0.163983 0.188144
Feature 10 -0.072835 0.004554 0.137737 0.170300
Feature 11 0.329457 0.023663 0.179561 0.141446
Feature 12 0.366667 0.151235 0.141446 0.219746
Feature 13 0.217842 0.005397 0.163901 -
Feature 14 0.288235 0.029469 0.105973 -
Feature 15 0.409091 0.008311 0.014444 -
Feature 16 0.343137 0.155172 0.105973 -
Feature 17 0.270492 0.269231 0.160693 -
Feature 18 0.203125 - 0.235832 -
Feature 19 - - 0.164412 -
Feature 20 - - 0.093699 -
Feature 21 - - -0.053386 -
Feature 22 - - 0.254810 -
Feature 23 - - -0.156301 -
Table 4: Features extracted from Slogan and Title datasets by PosData and MixedData. Values are subtracted by 0.5 to see values under 0.5 easily.
As shown in Table 4, several features from PosData and MixedData lacked the discriminative power and validity required for reliable analysis. In slogans, 3 out of 18 PosData features had DP scores below 0.1 (with 1 negative), and 8 out of 17 MixedData features were below 0.1. Similarly, in titles, 4 out of 23 PosData features were under 0.1 (with 3 negative), and 2 out of 13 MixedData features were under 0.1.
These results highlight that directly generating features from LLMs often produces features that lack sufficiently discriminative strength and reliability for robust data analysis, suggesting the need for improved methodologies.
  Category Bucket Feature 1 Feature 2 Feature 3 Feature 4 Feature 5 Feature 6 Feature 7 Feature 8 Feature 9 Average
  Slogan Bucket 1 0.3000 0.2969 0.1988 0.1791 0.1647 0.1446 0.1429 0.1114 0.0071 0.1644
Bucket 2 0.2963 0.2445 0.2442 0.2108 0.1828 0.1450 0.1050 0.0742 0.0126 0.1556
Bucket 3 0.2860 0.3000 0.2057 0.2034 0.2018 0.1715 0.1424 0.1149 0.1099 0.1878
Bucket 4 0.3000 0.2921 0.3072 0.3101 0.3097 0.3021 0.3239 0.2311 0.2810 0.3016
Bucket 5 0.3709 0.3239 0.3072 0.3000 0.2909 0.2810 0.2311 0.2921 0.3072 0.3016
Title Bucket 1 0.5000 0.4358 0.3438 0.1232 0.1000 0.0778 0.0625 0.0504 0.0493 0.1755
Bucket 2 0.3636 0.2935 0.2826 0.2692 0.1355 0.0997 0.0781 0.0576 0.0560 0.1685
Bucket 3 0.4379 0.4032 0.3704 0.3085 0.2953 0.1464 0.1268 0.0907 0.0897 0.2349
Bucket 4 0.4146 0.3699 0.3201 0.2995 0.1524 0.1506 0.1314 0.1259 0.1184 0.2182
Bucket 5 0.5000 0.4730 0.4545 0.4489 0.4389 0.4231 0.4213 0.3905 0.2822 0.4095
Table 5: DP scores for Slogan and Title Buckets. Values are subtracted by 0.5 to see values under 0.5 easily.
F.2 Discriminative Power of DSAI-Generated Features
To evaluate the reliability of features from DSAI, we organized them into five buckets along the prominence intensity spectrum. From each bucket, we randomly sampled 10 features to annotate the entire dataset with Conform/Not Conform labels when applicable, and marked Not-Applicable when features were inapplicable. For example, we marked "Avoid focusing solely on financial aspects when describing the business model" as Not-Applicable for data points that did not have any business model descriptions.
The analysis revealed that features in the lowest-prominence bucket (Bucket 1) exhibited lower discriminative power across both slogan and title datasets, while features in higher-prominence buckets demonstrated greater reliability (Table 5). This highlights the importance of implementing prominence-based thresholds to filter out less effective features. Notably, none of the DSAI-generated features had a DP score below zero, indicating that no feature mischaracterized negative data traits as positive. This result contrasts sharply with PosData and MixedData) which produced some features with negative DP scores.
These results validate that DSAI systematically filters out less impactful or unreliable features with the use of prominence-based thresholds, enabling more precise and reliable data annotation.
Appendix G Prominence Intensity and Occurrence Analysis
This section explores the relationship between feature frequency, prominence, and their impact on discriminative power. By analyzing feature occurrence and prominence distributions, we provide insights to help readers determine appropriate thresholds for these factors, aiding in the selection of features that optimize performance in distinguishing positive and negative traits within datasets.
(a) Average of Prominence Score per Frequency Bucket
(b) Data Coverage per Prominence Bucket
Figure 5: Comparison of Prominence scores and data coverage across frequency and prominence buckets.
G.1 Influence of Feature Occurrence on Discriminative Power
We analyzed the influence of feature occurrence frequency on discriminative power by grouping features into frequency buckets (e.g., <=10, 20-100, and >100) and evaluating mean/maximum prominence metrics (Figure 5(a))
Low-Frequency Features
Features with low frequencies (<= 10) demonstrated higher discriminative power, with mean prominence as high as 0.482 (some reaching even 1.0) in the slogan dataset. This suggests that low-frequency features effectively represent specific traits of positive data, though their rarity may limit generalization.
Medium-Frequency Features
Features with medium frequencies 20–100) displayed a balance between specificity and generality. They demonstrated consistent performance in discriminative tasks through moderate mean prominence values.
High-Frequency Features
Features with high frequency (> 100) were found to be more generic, showing low mean prominence values (0.087 in the slogan dataset for frequencies > 500). Though less discriminative, they provide valuable insights into the dataset’s baseline characteristics. Their high frequency makes them suitable for applications where specificity is less critical.
G.2 Influence of Feature Prominence on Discriminative Power
We also analyzed the distribution of features across different prominence score levels of by examining frequency accumulation within prominence buckets (Figure 5(b)). This analysis revealed how features were concentrated across low, moderate, and high prominence scores.
Slogan
The features were concentrated in the lower prominence range, with 61 features below 0.1 and 108 features below 0.2. This distribution suggests that while the majority of features exhibit limited discriminative power, a select subset of features with higher intensities plays a disproportionate role in capturing latent positive traits.
Title
An even more pronounced skew toward lower intensities was observed, with 11 features showing zero prominence and 62 features below 0.2. This distribution highlights that while low-prominence features comprise the majority of the dataset, they have limited effectiveness in distinguishing positive traits.
Appendix H Threshold Analysis on Slogans
In this section, we track how different prominence and frequency thresholds affect the recall of expert criteria, with particular attention to which criteria types are most resistant to threshold increases. In general, we observe progressive filtering of features as thresholds increase, which aligns with our criteria importance categorization. However, we also observed some exceptions diverging from their importance categorization. These findings illustrate how DSAI can identify discrepancies between theoretical feature importance and actual implementation patterns in the data.
H.1 Analysis of Expert-defined Criteria
We categorize expert criteria based on their importance to slogan effectiveness, ranging from critical to supplementary features. This categorization provides a framework for analyzing which features persist across different thresholds.
(a) Core Message Delivery features (Critical Importance)
These features are essential for effective communication of the brand’s core sales proposition and unique characteristics. They are considered critical because they directly influence a slogan’s ability to capture and deliver the brand’s main message to consumers.
  • 4. Convey the sales idea clearly and concisely
  • 8. Emphasize the brand’s unique traits or the benefits it provides
(b) Structure and Expression features (High Importance)
These features focus on clarity and readability, making a slogan easy to understand and leaving a lasting impression. They significantly impact audience engagement and retention, though they may not directly affect the message’s content.
  • 1. It should be direct and straightforward
  • 2. Keep it simple, but not overly simple
  • 5. Avoid misleading or exaggerated words
(c) Tone and Atmosphere features (Moderate Importance)
These features create an emotional connection with the audience through engaging and positive tone. While they enhance a slogan’s appeal, they are not as pivotal as core message delivery or structural clarity.
  • 3. It should have a pleasant tone
  • 6. It should be future-oriented
(d) Supplementary features (Lower Importance)
These features can enhance the slogan’s overall quality or effectiveness, but are not essential for a slogan’s primary function.
  • 9. Include the brand name in the slogan
  • 7. Clear positioning through comparison or closeness
H.2 Prominence Intensity Threshold Analysis
Table 6 displays the effect of different prominence thresholds on recall and data size.
  Recall Threshold (Prominence) Data Size
1 0.003 235
0.889 0.348 83
0.778 0.549 44
0.667 0.714 15
0.556 0.750 12
0.444 0.833 6
0.333 0.857 5
0 1 0
Table 6: Recall and Data Size Across Prominence Thresholds
Recall remains stable until reaching a relatively high threshold (0.348), indicating robust baseline coverage of our pipeline. The relationship between importance levels and threshold resilience shows an overall aligned pattern, though with some complexity: while lower-importance criteria are consistently filtered out early, criteria of moderate importance and above show varied retention patterns, with some excluded at mid-level thresholds and others persisting until the highest thresholds.
Figure 6: Dropped criterion as Prominence threshold increases
  1. 1.
    Early Exclusions (<0.6): 7. Clear positioning through comparison or closeness (lower importance) and 9. Include the brand name in the slogan‘ (lower importance) are excluded first, consistent with their supplementary nature.
  2. 2.
    Mid-level Exclusions (<0.8): General criteria such as 2. Keep it simple but not overly simple (high importance) and 4. Convey the sales idea clearly and concisely (critical importance) are excluded only at higher thresholds (0.721 and above), underscoring their broad applicability.
  3. 3.
    Late (No) Exclusions: Multiple features with moderate (6. Future-oriented) or above (e.g., 1. Directness (high importance), 8. Emphasize the unique traits and benefits (critical importance)) persist until the highest thresholds.
H.3 Frequency Threshold Analysis
Table 7 illustrates how recall changes as we filter features based on their frequency in the dataset.
  Recall Threshold (Frequency) Data Size
1 5 236
0.889 93 86
0.778 217 66
0.667 571 38
0.556 661 31
0.444 1,005 16
0.333 1,115 9
0.111 1,192 2
0 1,196 1
Table 7: Recall and Data Size Across Frequency Thresholds
Recall remains stable up to a frequency threshold of 93, indicating strong coverage of expert criteria even when considering only frequently observed patterns. The relationship between theoretical importance and threshold resilience reveals both expected alignments and interesting disparities. Critical features persist at high thresholds, and certain lower-importance features being filtered early. Some theoretically important features drop out earlier than expected, while certain lower-importance features demonstrate surprisingly high frequency in practice.
Figure 7: Dropped criterion as Frequency threshold increases.
  1. 1.
    Early Exclusions (<500): Notably, 5. Avoid misleading words (high importance) and 6. Future-oriented tone (moderate importance) drop out first. According to our theoretical categorization, these are not supplementary features, yet they appear less frequently in actual slogans than their theoretical significance would suggest.
  2. 2.
    Mid-level Exclusions (<1000): 3. Pleasant tone (moderate importance) and 9. Include brand name (lower importance) are filtered out at moderate thresholds, aligning with their moderate importance categorization.
  3. 3.
    Late Exclusions: Multiple core features like 4. Convey the sales idea clearly (critical importance) and 1. Direct and straightforward (high importance) has high frequency, confirming that these features are both theoretically critical and practically prevalent. 7. Clear positioning through comparison (lower importance) shows a notable deviation between theoretical and practical implementation.
Appendix I Threshold Analysis on Research Titles
Similar to our analysis on slogans (Appendix H), the results suggest that DSAI can effectively capture the nuanced reality of title construction, where practical implementation patterns may differ from theoretical guidelines.
I.1 Analysis of Expert-defined Criteria
We begin by categorizing expert-defined criteria based on their contribution to a title’s primary function: effective delivery of research content to readers.
Core Information Delivery features (Critical Importance)
These features are fundamental as they directly affect a title’s ability to enable readers quickly grasp the paper’s topic and contributions.
  • 3. The title needs to be informative and specific
  • 4. The title needs to be functional (with essential scientific "keywords")
  • 6. The title should include the main theme of the paper
Structural and Format features $\rightarrow$ Readability (High Importance)
These features optimize the title’s readability and clarity. While they do not directly affect content, they are crucial for successful information delivery.
  • 1. The title needs to be simple in terms of format
  • 2. The title needs to be direct
  • 5. The title should be concise and precise
  • 7. The title should not be too long or too short
Linguistic Expression features (Moderate Importance)
These features maintain academic professionalism while ensuring accessibility. They enhance the title’s effectiveness without being critical to its basic function.
  • 8. The title should avoid whimsical or amusing words
  • 9. The title should avoid non-standard abbreviations and unnecessary acronyms (or technical jargon)
Supplementary Guidelines (Lower Importance)
These guidelines can be advantageous in specific contexts but are not universally essential.
  • 10. Place of the study and sample size should be mentioned only if it adds to the scientific value of the title
  • 11. Important terms/keywords should be placed at the beginning of the title
  • 12. Descriptive titles are preferred to declarative or interrogative titles
I.2 Prominence Intensity Threshold Analysis
  Recall Threshold (Prominence) Data Size
0.833 0 199
0.750 0.692 40
0.667 0.778 29
0 1 0
Table 8: Recall and Data Size Across Prominence Thresholds
Table 8 demonstrates the impact of increasing the prominence threshold on the recall and data size.
Figure 8: Dropped criterion as Prominence threshold increases.
The majority of features are retained even at a high threshold of 0.692, suggesting that our methodology is robust in covering essential features. Analysis of the exclusion pattern demonstrates a general relationship between feature importance and retention, although some show divergence from their importance categorization:
  1. 1.
    Early Exclusions (<0.6): None.
  2. 2.
    Mid-Level Exclusions (<0.8): Style-related, moderate-importance features including 7. Not too long or short and 8. Avoid whimsical words and are excluded at moderate thresholds.
  3. 3.
    Late (No) Exclusions: Features of higher importance related to core message delivery and readability remain intact until the highest thresholds, underscoring their fundamental nature. Notable exceptions are 9. Avoid non-standard abbreviations and 12. Descriptive type, which are retained despite their mid to lower importance.
I.3 Frequency Threshold Analysis
We observe recall changes across varying ‘Frequency‘ thresholds.
  Recall Threshold (Frequency) Data Size
0.833 1 284
0.750 791 44
0.667 969 32
0.583 1,233 26
0.500 1,294 20
0.417 1,296 19
0.333 1,326 13
0.250 1,333 11
0.167 1,344 4
0.083 1,345 1
Table 9: Recall and Data Size Across Frequency Thresholds
Recall remains stable until a threshold of 791. The relationship between theoretical importance and retention patterns again reveals some unexpected deviations: a high-importance feature drops out early, while other features of similar importance persist until high thresholds. Additionally, while some supplementary features are not recalled at all, the recalled one (12. Descriptive type) shows remarkably strong retention, suggesting that practical title construction may prioritize certain features differently from theoretical guidelines.
Figure 9: Dropped criterion as Frequency threshold increases
  1. 1.
    Early Exclusions (<500): None.
  2. 2.
    Mid-Level Exclusions (<1000): Features ranging from moderate (9. Avoid non-standard abbreviations) to critical importance (6. Include the main theme) are excluded at this stage.
  3. 3.
    Late Exclusions: Structural features like 1. Simple format (high importance), 5. Concise and precise (high importance) persist until the highest thresholds, aligning with their theoretical importance. Interestingly, 12. Descriptive type (low importance) shows the highest retention despite its low theoretical importance.
Appendix J Reliability of Pipeline Operations
In this section, we assess the reliability of our methodology by examining potential errors that could compromise the data-groundedness of the generated features. We specifically focus on cases where incorrect mapping between datapoints and features undermines the accuracy of the feature extraction process.
#1 Perspective Generation
This stage is conceptually error-free by design, as there is no strict requirement for direct correspondence between the generated perspectives and the dataset. While the dataset provides heuristic guidance, the perspectives can be freely generated without compromising the methodology’s integritiy. Therefore, the concept of errors is not applicable to this stage.
#2 Perspective Value Matching Reliability check
Errors can occur when incorrect values are generated, resulting in improper mapping between datapoints and the corresponding perspective-value pairs. To evaluate this, we performed a consistency check on the generated values using LLMs, examining at least three distinct sentences per perspective. The consistency check yielded high accuracy rates exceeding 98% for both research titles and slogans. Specifically, the error rates were 0.042 for slogans (11 out of 258), 0.071 for titles (29 out of 405), and 0.053 for the MIND dataset (15 out of 279). The prompt we used for consistency check can be found in Figure 28.
#3 Value Clustering Reliability check
Errors in this stage occur when cluster labels are incorrectly generated, causing mis-mapping during the conversion from perspective-value to perspective-label forms. This mis-mapping can distort the intended analysis. We conducted the same consistency check as in stage #2, substituting values with labels as input. The accuracy rates exceeded 94% for both titles and slogans.
#4 Verbalization Reliability check
Errors can also arise during the verbalization of perspective-label pairs. To evaluate this, we selected 20 features from each feature bucket generated in Section F.2. We assessed each feature according to two criteria: (1) whether the sentence captures the combined meaning of the perspective and label, and (2) whether the sentence correctly applies directional score by rewriting it in a direct form if $P(\text{positive}|\text{conform})>0.5$, or in an avoidant form if $P(\text{positive}|\text{conform})\leq 0.5$. This evaluation used a detailed prompt requiring True/False responses for each criterion, with returned results in JSON format. The verbalization reliability check demonstrated high accuracy exceeding 98% for both titles and slogans, confirming that the generated sentences faithfully represent the intended meaning. The prompt we used for reliability check can be found in Figure 29.
#5 Prominence-based Selection
This stage is inherently error-free as it operates deterministically through mathematical calculations without involving LLMs.
Appendix K DSAI-Generated Top/Bottom Prominence Features of Expert-Driven Annotation Dataset
As discussed in Section 5.1, the features generated using the slogan and title datasets demonstrate high recall values when compared to expert-defined requirements. This holds true even when applying a high threshold for prominence, indicating that the generated features effectively capture the key characteristics outlined by experts. However, as highlighted in Section 5.2, while features with high prominence exhibit strong discriminative power and reliability, those with lower prominence tend to have relatively lower reliability.
To provide further insights, we present the top and bottom 20 features for each dataset along with their prominence scores and frequencies. The results for the slogan dataset are detailed in Table 11, and those for the title dataset are outlined in Table 12.
Appendix L DSAI-Generated Top/Bottom Prominence Features of Industry Dataset
L.1 Prominence Distribution
Figure 10: Distribution of each dataset based on prominence scores.
The following analysis evaluates the prominence score distribution of criteria across three datasets: Reddit, MIND, and SPAM. These scores indicate how reliably each feature identifies the latent characteristics within the datasets.
MIND
The MIND dataset shows a highly concentrated prominence distribution, with most scores reaching 1.0. Few features fall below 0.85, reflecting strong alignment between the generated features and the structured nature of the dataset. These consistently high scores validate the methodology’s effectiveness in extracting meaningful patterns from structured data with CTR rates.
Spam
The Spam detection dataset exhibits a broader prominence distribution than MIND. Still, a large portion are above 0.9, indicating that the methodology is generally effective for SPAM. However, a small subset scoring below 0.5 indicates features that may be less suitable for classification purposes, suggesting the need for careful filtering of specific features for robust classification.
Reddit
The Reddit dataset displays a broad and uneven distribution of prominence scores, peaking at 0.67 with a gradual decline toward lower scores. A significant number of features cluster below 0.4, with many falling below 0.1. This distribution reflects the unstructured and highly diverse nature of Reddit content, making consistent pattern identification challenging. The broad range of topics and content variability suggests that individual features may be highly specific to certain data subsets while falling to generalize across others.
L.2 General vs Specific Features
  Perspective Label Type Frequency
                                 General Features
  reasoning Lack of Logical Reasoning NEGATIVE 1,038
fictional_reference Absent NEGATIVE 1,033
gender_roles no_reference POSITIVE 1,014
archaic_reference no_reference NEGATIVE 1,004
euphemistic_language euphemism_absent NEGATIVE 986
historical_reflection Absent NEGATIVE 966
partisan_tone absent POSITIVE 964
legal_context Absence POSITIVE 962
objective_evaluation Subjective NEGATIVE 950
validity subjective NEGATIVE 926
scientific_reference No Scientific Reference NEGATIVE 918
profanity Profanity Absent POSITIVE 910
                                 Specific Features
  specific_undertone Sarcastic NEGATIVE 116
theme_recognition Justice and Morality NEGATIVE 116
cultural_sensitivity high_sensitivity POSITIVE 114
efficacy Ineffectiveness POSITIVE 114
comparison Indirect Comparison NEGATIVE 114
pragmatic Criticism NEGATIVE 110
audience_engagement Indirect Engagement POSITIVE 110
irony Intensity_Irony NEGATIVE 110
conflict Presence POSITIVE 104
demographic_target Youth NEGATIVE 102
emotionally_charged Emotionally Intense NEGATIVE 102
length two_digit_high POSITIVE 102
Table 10: Features and their corresponding labels, types, and frequency.
DSAI-generated features capture both general characteristics shared across the dataset and highly specific features. As shown in Table 10, some features represent broad, overarching traits that are present across a significant portion of the dataset, with high frequencies (900+ instances). These include common patterns such as "lack of logical reasoning" or "absence of historical reflection," which are applicable across diverse contexts.
In contrast, more specific features capture nuanced and detailed attributes that apply to smaller data subsets, appearing in fewer samples (1̃00 instances), such as "sarcastic undertone" or "emotionally intense". These characteristics necessitate fine-grained, data-driven analysis and are typically challenging for LLMs to identify due to their limited presence in pre-training data and the source dataset. This capability to extract such specific patterns beyond general trends demonstrate DSAI’s unique strength over direct feature extraction from LLMs.
The top and bottom 20 prominence fearures for MIND, SPAM and Reddit dataset are provided in Table 13, 14, 15, respectively.
  Requirement Prominence Frequency
                         Top Requirements
  The advertising tone should convey a focus on quality, using optimistic and aspirational language. 0.8571 14
The sentence should convey an optimistic advertising tone that encourages engagement. 0.8333 108
The sentence should incorporate cultural references that align with consumptive themes. 0.8333 12
The sentence should employ indirect methods to engage the audience effectively. 0.7857 28
Ensure the sentence contains a component that emotionally appeals to the reader. 0.7831 83
The sentence should effectively incorporate figurative language to enhance meaning. 0.7556 90
The sentence should convey a high level of prominence using strong and emotive language. 0.7215 79
Ensure the sentence conveys its message concisely, avoiding unnecessary words or lengthy expressions. 0.7143 14
The sentence should clearly present a promise of quality, assuring trust and excellence. 0.6406 217
The sentence should employ direct engagement techniques to capture the audience’s attention. 0.6250 128
The sentence should avoid losing its identity by being overly generic or broad in purpose. 0.6190 21
Ensure grammatical accuracy throughout the sentence. 0.6154 26
Avoid using the ampersand (&) in formal writing or titles. 0.5882 34
The sentence should clearly define and communicate a distinct value proposition. 0.5870 184
Ensure the sentence effectively conveys its intended message. 0.5789 19
The sentence should convey a strong and clear emotional tone. 0.5709 275
The sentence should provoke thought and engage the audience in a meaningful way. 0.5556 18
The sentence should avoid presenting the main idea in a way that misaligns with the intended topic or domain. 0.5556 18
The sentence should use persuasive language to effectively encourage action or belief. 0.5528 407
Ensure the sentence includes the mention of a relevant company or organization. 0.5493 213
                         Bottom Requirements
  The sentence should include a clear and compelling call to action. 0.0294 1,088
The sentence should fully and effectively communicate its intended message. 0.0267 1,159
The sentence should avoid merely providing information without an intended action or emotion. 0.0250 1,122
Ensure the sentence includes references to cultural significance. 0.0248 1,009
Avoid using imperative or overly complex grammatical structures in titles. 0.0244 41
Ensure the use of inclusive language in the sentence. 0.0225 1,109
Avoid sentences that lack necessary cultural references. 0.0224 1,115
The sentence should not be overly specific, limiting its relevance to a narrow audience. 0.0205 537
The sentence should maintain a general level of specificity to appeal to a wide audience. 0.0166 661
Ensure that sentences maintain grammatical correctness. 0.0162 1,171
The sentence should focus on delivering specific and detailed information. 0.0148 741
Avoid using general language and focus on providing specific details in the sentence. 0.0131 458
The sentence should be concise, conveying its message briefly and efficiently. 0.0110 1,185
The sentence should maintain clear and direct language. 0.0095 1,157
The sentence should avoid overly focusing on the business aspect at the expense of other dimensions. 0.0070 572
The sentence should explicitly mention the presence of a specific service being offered. 0.0048 1,043
The sentence should effectively convey clear and useful information. 0.0043 1,163
The sentence should maintain simplicity in structure and composition. 0.0034 ,192
The sentence should be composed in a single language for consistency. 0.0033 1,196
Avoid emotionally neutral language that fails to connect with readers on an emotional level. 0.0032 315
Table 11: Top and Bottom Requirements of Slogan Dataset based on Prominence.
  Requirement Prominence Frequency
                         Top Requirements
  Avoid using list structures that compromise grammatical integrity and clarity. 1.0000 11
Avoid vague and nonspecific language in the sentence. 1.0000 11
Avoid the use of pronominal subjects in sentences. 1.0000 12
Ensure the sentence provides a clearer focus on the topic. 1.0000 16
Avoid using an informal narrative style in the sentence. 1.0000 33
Ensure the sentence maintains a strong technical focus without shifting to a non-technical direction. 1.0000 41
Ensure the inclusion of phrasal nouns in the sentence where appropriate. 0.9333 30
The sentence should clearly convey its purpose or intent. 0.9259 27
Ensure the sentence addresses the topic clearly without any ambiguity. 0.8854 192
Ensure the sentence clearly specifies its domain of application. 0.8750 16
The sentence should incorporate technical adjectives to enhance precision and clarity in describing nouns or pronouns. 0.8333 12
The sentence should clearly highlight learning methods as the main topic. 0.8182 11
Avoid using interrogative sentences. 0.8182 33
Ensure complete grammatical structures in the sentence. 0.8095 21
Avoid excessive use of concrete nouns in the sentence. 0.7857 28
Avoid overly complex or frequent use of interrogative structures in the content. 0.7838 37
The sentence should avoid being overly verbose and aim for syntactic compression. 0.7739 115
Ensure sentences are concise and avoid unnecessary length. 0.7500 40
The sentence should clearly use either passive or active voice to avoid ambiguity. 0.7333 30
Avoid focusing on concrete, tangible items in the sentence. 0.7000 40
                         Bottom Requirements
  Ensure modal verbs are absent in the sentence if not necessary for conveying meaning. 0.0840 917
Limit the length of main noun phrases to enhance readability. 0.0833 96
The main noun phrase in the sentence should ideally consist of two words. 0.0833 144
Avoid using sentence fragments to ensure complete and meaningful sentences. 0.0805 174
Avoid using definite articles when introducing new or less familiar concepts. 0.0748 147
The sentence should effectively use past tense to communicate events or actions that have already occurred. 0.0667 15
The sentence should embrace high complexity in its structure and vocabulary. 0.0581 172
Avoid sentences with overly simplistic structures or vocabulary. 0.0575 435
Ensure sentences utilize participles effectively to enhance clarity and detail. 0.0529 473
Ensure the sentence includes infinitive verb forms for clarity and action orientation. 0.0526 57
Avoid reliance on non-verbal tense constructs in the sentence. 0.0500 40
The sentence should clearly identify and be applicable to the educational domain. 0.0476 21
The sentence should maintain a neutral tone by balancing between passive and active voice. 0.0476 84
Limit the use of multiple verbs in a single sentence. 0.0455 88
Avoid sentences that inappropriately mix technical terms outside the humanities and social sciences domain. 0.0400 25
Limit the use of multiple adjectives in a sentence to maintain clarity and precision. 0.0357 56
The sentence should clearly articulate its relationship to the field of Health and Medicine. 0.0323 31
Ensure the sentence includes well-chosen adjectives to enhance description and clarity. 0.0121 330
Ensure the presence of a clear tense in the sentence. 0.0076 264
Ensure the presence of finite verbs to convey clear action or timing in the sentence. 0.0039 510
Table 12: Top and Bottom Requirements of Title Dataset based on Prominence.
  Requirement Prominence Frequency
                         Top Requirements
  The sentence should effectively convey a sense of distress through emotional language. 1.0000 12
Avoid language that implies assistance or support for actions that should be independent. 0.8462 13
Avoid statements that may have an unintended negative impact on the audience. 0.7895 19
Avoid overloading the sentence with complex or unnecessary details related to technology and cybersecurity. 0.7241 29
The sentence should effectively convey a sense of frustration through emotional language. 0.7143 14
The sentence should aim to surprise or alarm the audience to elicit a strong reaction. 0.6667 24
Avoid using language that only implies future outcomes without clear details. 0.6667 18
Avoid using domain-specific language related to the military and defense. 0.6364 11
The sentence should convey information of lesser importance effectively. 0.6364 11
The sentence should avoid focusing solely on technology and innovation when identifying the audience. 0.6364 11
The sentence should clearly specify the legal or criminal aspects of a particular field. 0.6000 35
Avoid using language that fails to motivate or inspire the audience. 0.5625 32
Ensure the main subject of the sentence pertains to sports. 0.5556 18
The sentence should clearly reference past legal cases using appropriate legal terminology. 0.5385 26
Avoid using quality-based adjectives that may imply subjective judgment or bias. 0.5000 24
The sentence should not overly highlight negative issues that could alarm the audience. 0.5000 12
Avoid referencing military actions, personnel, or events in the content. 0.4737 38
The sentence should avoid focusing solely on social and lifestyle aspects within its domain. 0.4737 19
The sentence should avoid using action verbs that emphasize creation-related actions. 0.4545 11
Avoid using ambiguous modal verbs when expressing plans or intentions. 0.4444 18
                         Bottom Requirements
  Ensure the inclusion of specific numerical data or statistics in the sentence. 0.0078 766
Avoid overly clear or explicit language that might be inappropriate or too revealing in certain contexts. 0.0073 961
Ensure the sentence does not lack attention to potential contradictions. 0.0071 989
The sentence should exclude references to social media platforms. 0.0062 966
The sentence should effectively communicate its message without relying on figurative language. 0.0059 678
The sentence should avoid the use of separators such as dashes, colons, or slashes. 0.0054 742
Ensure the inclusion of sensory details in the sentence to enhance vividness. 0.0051 973
The sentence should clearly provide instruction or guidance to the reader. 0.0051 979
The sentence should present concrete, specific ideas and details. 0.0042 946
Ensure that the sentence is clear and needs no further clarification. 0.0042 952
Ensure that scientific theories or principles are present in the sentence. 0.0041 976
The sentence should avoid the use of hashtags. 0.0040 994
The sentence should provide sufficient context to be understood independently. 0.0035 865
Avoid or limit the use of figurative language in the sentence. 0.0032 313
Avoid sentences that lack necessary negations to clarify intended meaning. 0.0031 977
Avoid using ambiguous or unsuitable speech acts in sentence construction. 0.0030 989
The sentence should avoid hypothetical scenarios to maintain factual clarity. 0.0022 898
The sentence should avoid content that lacks relevance or significance, especially for a news context. 0.0020 980
The sentence should clearly identify the language being used, ensuring linguistic features are consistent with the specified language. 0.0020 998
The sentence should be free from unnecessary repetition, ensuring clarity and conciseness. 0.0010 985
Table 13: Top and Bottom Requirements of MIND dataset based on Prominence.
  Requirement Prominence Frequency
                         Top Requirements
  Avoid using list structures that compromise grammatical integrity and clarity. 1.0000 11
Avoid vague and nonspecific language in the sentence. 1.0000 11
Avoid the use of pronominal subjects in sentences. 1.0000 12
Ensure the sentence provides a clearer focus on the topic. 1.0000 16
Avoid using an informal narrative style in the sentence. 1.0000 33
Ensure the sentence maintains a strong technical focus without shifting to a non-technical direction. 1.0000 41
Ensure the inclusion of phrasal nouns in the sentence where appropriate. 0.9333 30
The sentence should clearly convey its purpose or intent. 0.9259 27
Ensure the sentence addresses the topic clearly without any ambiguity. 0.8854 192
Ensure the sentence clearly specifies its domain of application. 0.8750 16
The sentence should incorporate technical adjectives to enhance precision and clarity in describing nouns or pronouns. 0.8333 12
The sentence should clearly highlight learning methods as the main topic. 0.8182 11
Avoid using interrogative sentences. 0.8182 33
Ensure complete grammatical structures in the sentence. 0.8095 21
Avoid excessive use of concrete nouns in the sentence. 0.7857 28
Avoid overly complex or frequent use of interrogative structures in the content. 0.7838 37
The sentence should avoid being overly verbose and aim for syntactic compression. 0.7739 115
Ensure sentences are concise and avoid unnecessary length. 0.7500 40
The sentence should clearly use either passive or active voice to avoid ambiguity. 0.7333 30
Avoid focusing on concrete, tangible items in the sentence. 0.7000 40
                         Bottom Requirements
  Ensure modal verbs are absent in the sentence if not necessary for conveying meaning. 0.0840 917
Limit the length of main noun phrases to enhance readability. 0.0833 96
The main noun phrase in the sentence should ideally consist of two words. 0.0833 144
Avoid using sentence fragments to ensure complete and meaningful sentences. 0.0805 174
Avoid using definite articles when introducing new or less familiar concepts. 0.0748 147
The sentence should effectively use past tense to communicate events or actions that have already occurred. 0.0667 15
The sentence should embrace high complexity in its structure and vocabulary. 0.0581 172
Avoid sentences with overly simplistic structures or vocabulary. 0.0575 435
Ensure sentences utilize participles effectively to enhance clarity and detail. 0.0529 473
Ensure the sentence includes infinitive verb forms for clarity and action orientation. 0.0526 57
Avoid reliance on non-verbal tense constructs in the sentence. 0.0500 40
The sentence should clearly identify and be applicable to the educational domain. 0.0476 21
The sentence should maintain a neutral tone by balancing between passive and active voice. 0.0476 84
Limit the use of multiple verbs in a single sentence. 0.0455 88
Avoid sentences that inappropriately mix technical terms outside the humanities and social sciences domain. 0.0400 25
Limit the use of multiple adjectives in a sentence to maintain clarity and precision. 0.0357 56
The sentence should clearly articulate its relationship to the field of Health and Medicine. 0.0323 31
Ensure the sentence includes well-chosen adjectives to enhance description and clarity. 0.0121 330
Ensure the presence of a clear tense in the sentence. 0.0076 264
Ensure the presence of finite verbs to convey clear action or timing in the sentence. 0.0039 510
Table 14: Top and Bottom Requirements of Spam Detection Dataset based on Prominence.
  Requirement Prominence Frequency
                         Top Requirements
  The sentence should avoid unnecessary criticism and focus on constructive feedback. 0.6727 110
The sentence should avoid using strong partisan language or political bias. 0.6129 124
The sentence should include elements that evoke a surreal or dream-like quality. 0.6104 154
The sentence should avoid being perceived solely as a review. 0.6082 194
The sentence should effectively convey a positive emotion. 0.5699 186
The sentence should effectively convey information and facilitate informative sharing with the audience. 0.5333 120
Avoid using a critical tone in the sentence. 0.5313 128
The sentence should convey a positive sentiment. 0.5294 136
Avoid language that inaccurately places the sentence within the politics and media domain. 0.5238 126
Ensure the sentence does not include language indicating complaints or grievances. 0.5130 230
The sentence should align with public opinion and mainstream consensus. 0.5062 162
The sentence should address topics in a non-sensitive manner, avoiding contentious or inflammatory language. 0.5044 226
Avoid displaying a dominant interpersonal stance in relational dynamics. 0.4907 161
The sentence should convey a positive emotional tone. 0.4815 108
The sentence should convey information with high intensity and detail to maximize audience knowledge gain. 0.4737 190
The content should maintain an objective, unbiased perspective throughout. 0.4729 129
The sentence should use words with optimistic or positive connotations. 0.4516 124
The language used should maintain a neutral tone, avoiding extreme politeness or rudeness. 0.4476 210
Avoid using circular or flawed reasoning in arguments. 0.4468 188
The sentence should avoid rhetorical devices and maintain a straightforward approach. 0.4430 237
                         Bottom Requirements
  The content should maintain a moderate level of complexity. 0.0116 518
Avoid non-impressionistic language and incorporate more vivid or subjective descriptions. 0.0110 182
The sentence should avoid legal language or references. 0.0104 962
The sentence should avoid suggesting a lack of consensus or collective agreement. 0.0099 202
The sentence should appropriately reference hierarchical authority to establish credibility or significance. 0.0093 214
Avoid using American regional dialects or cultural references in the sentence. 0.0090 222
The sentence should avoid depicting any specific gender roles or influences. 0.0081 867
Ensure the sentence includes relevant jargon where necessary to convey expertise and precision. 0.0080 754
The sentence should avoid overly intense expressions of real-world relevance. 0.0062 1,129
Ensure the sentence length is concise, ideally with a word count in the single digits. 0.0061 330
The sentence should avoid making explicit factual assertions without sufficient evidence or context. 0.0060 668
The sentence should be inclusive and cater to a broad demographic audience. 0.0056 354
The sentence should clearly demonstrate an understanding of various sentence types and structures. 0.0053 1,126
The sentence should not restrict or misrepresent opinion sharing. 0.0047 426
The sentence should be intellectually demanding but still accessible at a medium level of complexity. 0.0045 446
Avoid using exclusive language or sentiments in the sentence. 0.0043 235
Minimize the use of overly dramatic or theatrical language in the sentence. 0.0027 375
The sentence should maintain a concise length, ideally within a low two-digit word count. 0.0026 389
The sentence should intentionally exclude hashtags. 0.0025 1,191
Ensure language identification and classification are accurate within the sentence. 0.0017 1,194
Table 15: Top and Bottom Requirements of Reddit dataset based on Prominence.
Figure 11: Expert-Defined Criteria
Figure 12: LLM Annotation prompt
Figure 13: Prompts for direct LLM feature generation.
Figure 14: NoData Result for slogan
Figure 15: PosData and FlippedPosData Results for slogan
Figure 16: MixedData and FlippedMixedData Results for slogan
Figure 17: NoContext Result for slogan
Figure 18: NoData Result for title
Figure 19: PosData Result for title
Figure 20: FlippedPosData Result for title
Figure 21: MixedData and FlippedMixedData Results for title
Figure 22: NoContext Result for title
Figure 23: #1 Perspective Generation Prompt and Outputs
Figure 24: #2 Perspective Value Matching Prompt and Outputs
Figure 25: #3 Clustering Prompt
Figure 26: #3 Clustering Outputs
Figure 27: #4 Verbalization Prompt and Outpus
Figure 28: #2 Perspective Value Matching Reliability Test Prompt
Figure 29: #4 Verbalization Reliability Test Prompt